OpenStudy (anonymous):

Find the slope of the tangent line to the curve 2x^2+4xy+2y^3=114 at the point (-7,4)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (triciaal):

do you know how to take the derivative of a function?

OpenStudy (triciaal):

when x = 0 y = 2Y^3 = 114 so y^3 = 57 y = 57^1/3 = 3.84 when y = 0 2x^2 = 114 x^2 = 57 x = + and - 57^1/2 = 7.55

OpenStudy (triciaal):

split the original into an x function and a y function

OpenStudy (triciaal):

@jim_thompson5910 what is the correct way to do this problem?

jimthompson5910 (jim_thompson5910):

you have to use implicit differentiation here

jimthompson5910 (jim_thompson5910):

here's a page on implicit differentiation http://www.sosmath.com/calculus/diff/der05/der05.html you basically derive both sides with respect to x

OpenStudy (triciaal):

so it's x and y as a function of x ? thanks

jimthompson5910 (jim_thompson5910):

yes that's correct, y = f(x)

OpenStudy (triciaal):

thanks for all your help

OpenStudy (anonymous):

i did implicit differentiation and got dy/dx= (4x-y-1)/x+9y^2 and it said its wrong, what did i do wrong?

jimthompson5910 (jim_thompson5910):